A student makes a $10.50 purchase at the bookstore with a $20.00 bill. The store has no bills and gives change in quarters and fifty-cent pieces. There are 30 coins in all. How many of each kind are there?
Answer by mananth(16946)

He has $9.50 returned by the store
There are 30 coins in quarters and fifties
Let the quarter be x
fifties will be 30-x
25x+50(30-x)=950
25x+1500-50x=950
-25x= -550
x=-550/-25
x=22
Number of quaters = 22
Number of fifties =8